Friday, September 22, 2017

9/22/17

TGM alarm maxed out first in pump enclosure then in both pump and reactor enclosures, with minimum readings at the scrubber outlet and system enclosure.

Found a crack pipe clamp on the vacuum line going to the pump.  Replaced this connecting pipe clamp.

Replaced the pump oil.

Wednesday, June 21, 2017

AsH3 cylinder was replaced. 
     Lot # 1987306991Q7  SV020731
     Net Weight: 1.98 lbs or 898.11 g

Old cylinder is being stored in back of MOCVD enclosure.

Wednesday, May 10, 2017

4-20-2017

Charcoal in scrubber was changed.  This was done to help prevent readings at the scrubber outlet when using PH3.

5-5-2017

A finer grade of charcoal was added to the charcoal in the scrubber.
5-10-2017

TGM pump stopped working during the night.  A fastener screw was sheared off and I was unable to adjust the flow at any of the points.  All points were reading low flow.  The pump was replaced with a new one, which is working fine. - NS

Friday, December 2, 2016

Tuesday, July 26, 2016

MOCVD maintenance work: 8-27-2015 ~ 7-7-2016


Historical records

8-27-2015 ~ 7-7-2016

Please refer to the original MOCVD System log book for more details.
--------------------------------------------------------------------------------------------------------------------------

-----------------------------
8-31-2015 ~ 9-1-2015
Deep cleaning of the vacuum pump and the oil trap:
clean the inside chamber that holds the pump oil;
clean the inside wall of the oil trap;
change vacuum pump oil;
change dust filter of the vacuum pump;
change to a new mist filter for the oil trap.
- NZ

-----------------------------
9-11-2015
Noticed that H2O sensors don't work
- PA

-----------------------------
9-25-2015
When brought the H2 out of idling, H2 purifier temperature went up to as high as 435 degC, at the cell evacuation step.

Reason: H2 flow rate=390 sccm and bleed valve was all the way open, thus H2 was evacuated at a very high rate. Bad for the cell.

Solution: Keep the bleed valve at 5 sccm bleeding rate until H2 flow rate reading (of the H2 MFC on the MOCVD panel) starts to drop from setting or stays at its minimum.
- NZ

-----------------------------
9-27-2015
Maybe a leak around AsGPV11.
H2 line never fully evacuates when As on/ purge/ H2 flush on.
Also As flow never drips below ~0.22 sccm.

Leak somewhere in H2 line on AsH3/ N-dope side.
- PA

-----------------------------
10-1-2015
Trouble with solenoid valves (previous day) (air leak). Removed part of segment 2 ( cable #52).

- PA

-----------------------------
10-4-2015
Changed pump oil, replace oil purifier filters.

Evacuated AsH3 line back to AGMV1;
Evacuated PH3 line back to PGMV1;
Evacuated TMAl line (twice, filled with N2, evacuated.)
Evacuated TMIn line (twice, filled with N2, evacuated.)
Need to refill lines before use.
- PA

-----------------------------
10-5-2015
Re-tuned heater.
- NS

-----------------------------
10-9-2015
Evacuated As and P lines.
Found/ Fixed wire going to 53: PGSV1.
- NS

-----------------------------
10-12-2015
Purging AsH3 line;
Set off TGM through high- pressure side. Not sure why.
Pump/Reactor/Scrubber cabinets.
- PA

-----------------------------
10-15-2015
AGMFC1 was removed and we opened it, and unfortunately, the sensor part of the MFC was made of micro-wired connecting to the electronic circuit and easy to break, thus we broke the micro-wireds and would try to fix it.

Label: AGMFC1
Make: MKS
Serial #: 021766633
Model #: M100B01311CS1BV
Range: 0~10 sccm

Put an old MFC on to replace AGMFC1
Range: 0~10sccm
Make: Vacuum General- Ultra-Flo
Serial #: 515810588
Model #: UC2-11S01

Dismount the H2 monitor from the wall, and connect it to a voltmeter. Test it with a small H2 flow and read the voltmeter.
Baseline: in air, min reading=0.85 volt.
When detects H2, reading increases, may go up to 20 volt or more.
- NZ

-----------------------------
10-16-2015
Check connections around the replaced MFC and manifold for leaks by filling the lines with H2.

Filled and evacuated the arsine and phosphine line for several times with H2 back to the AGMV1 and PGMV1.
-NS

-----------------------------
10-19-2015
Filled both arsine and phosphine lines with their respective source gas. Set regulator to 20 psi on both.
Put new chem. sensitive paper near the cylinders. Let both lines run for ~25 min.
- NS

-----------------------------
10-19-2015
Ran arsine line at different values to test how responsive and steady the MFC is.

Flow       Steadiness
(sccm)  
2             Good, no fluctuation
4             Good, no fluctuation
6             Good, no fluctuation
8             Good, no fluctuation
10           Good, no fluctuation
- NS
-----------------------------
10-20-2015
Created a Mo cover/holder to go on top of the susceptor.
- NS

-----------------------------
10-21-2015
Baked susceptor with Mo for ~20 min at 750 degC.
Mo became tinted brown in color.

Cleaned CARBOLITE furnace.
- NS

-----------------------------
11-8-2015
Reset TGM counter, need to change tape in the next few days.
- PA

-----------------------------
11-11-2015
Change TGM tape.
- PA

-----------------------------
12-14-2015
Flushed TMIn lines to weigh the bubbler.
TMIn bubbler was disconnected and placed in a container.
- NS

-----------------------------
12-16-2015
1939.7 g solution TMIn, S21123
Initial mass was 1944.6 g
- PA

-----------------------------
12-16-2015
Reattached and flushed TMIn source.
- NS

-----------------------------
12-21-2015
Brought H2 purifier out of idle, set H2 manifold to 300 and noticed purifier went to 425 degC.
- NS

-----------------------------
12-22-2015
New susceptor came in. 20 in total.
All 20 are individually wrapped and in a container labeled "New susceptor".

-----------------------------
1-13-2016
Deposition on reactor walls reached higher than usual.
- NS

-----------------------------
1-15-2016
Bleed valve will not fully close, and the bleed knob just keeps spinning clock-wise.
- NS

-----------------------------
1-21-2016
Fully opened and cleaned the reactor chamber.
Removed and cleaned bellows, thermal couple and rod.
Put everything back together and raised rod height to 6cm.
- NS

-----------------------------
1-25-2016
TMGa bubbler
1583.6 g P2292-21
- PA

-----------------------------
2-3-2016
H2 purifier got to 422 degC, bleed flow wasn't correct.
- NS

-----------------------------
2-10-2016
Replace N2 cylinder. New cylinder is at 2550 psi.
- NS

-----------------------------
2-15-2016
Replaced MOCVD susceptor rod.
-NS

-----------------------------
2-16-2016
Remove bellows.
KF40 adapters on chamber and rotary drive feed-through.
- PA

-----------------------------
2-17-2016
HP3 bleed valve is leaking.
Valve was replaced.
- NS and PA

-----------------------------
2-18-2016
HP3 leaked again. Replaced and greased.
HMV5 adjusted.
HP3 valve leaks when turned and closed all the way.
- PA

-----------------------------
2-22-2016
Flushed In line back to the bubbler for 10 times.
- NS

-----------------------------
2-23-2016
Flush TMIn for 20 times.
- PA

-----------------------------
2-24-2016
Flushing TMIn line for 5 times.
- PA

-----------------------------
2-24-2016
Added ball-bearing to thermal couple.
- NS

-----------------------------
2-24-2016
After growth, flushed TMIn for 5 times.
- NS

-----------------------------
2-25-2016
Changed back to flexible tubing to connect chamber.
- NS

-----------------------------
2-26-2016
After growth, flushed TMIn line for 5 times.
- NS

-----------------------------
2-29-2016
Change reactor o-ring.
- PA

-----------------------------
2-29-2016
After growth, flushed TMIn line for 5 times.
- NS

-----------------------------
3-1-2016
There were two gaskets on inlet line to reactor, removed, leak rate improved.
- PA

-----------------------------
3-2-2016
Leak rate is poor with flexible inlet.
Switch ed back to rigid inlet. Vacuum good.
- PA

-----------------------------
3-3-2016
Adjusted thermo-couple location to hopefully improve reading.
- NS

-----------------------------
3-4-2016
Chamber pressure ~342 Torr in the morning.
- PA

-----------------------------
3-4-2016
After growth, flush the TMIn line for 5 times using key.
- NS

-----------------------------
3-7-2016
Full purge of TMIn bubbler line with H2 for 8 times.
- PA

-----------------------------
3-7-2016
After growth, flushed TMIn line for 5 times.
- NS

After another growth, flushed TMIn bubbler line ~10 times.
- NS

-----------------------------
3-8-2016
Changed TMIn bath to 10 degC.
Changed PR2 to 10 psig (was ~18 psig).
- PA


-----------------------------
3-9-2016
Reset TGM tape counter to 120 days without changing the tape (4 days left).
- PA

-----------------------------
3-11-2016
Removed ALMFC1
MKS
S/N 021810488
0~100 sccm

Installed new (used) ALMFC1
Ultra-Flo
S/N 515760588
0~100 sccm

ALMFC1 unable to control flow at high flow setting (100 sccm) when switching pressure on.
- PA

-----------------------------
3-11-2016
Flushed Al lines 20 times with N2/ H2.
- NS

-----------------------------
3-14-2016
Test vacuum by plugging inlet and exhaust lines seems there is a very small leak somewhere in or around the manifold.
- NS

-----------------------------
3-15-2016
Flushed AsH3 line back to cylinder 5 times;
Flushed TMIn line back to bubbler 10 times;
Flushed Al line back to bubbler 10 times;
Flushed PH3 line back to cylinder 5 times.

Reactor cabinet alarm went off during first cycle.
Flush TMGa line back to bubbler 5 times.
- NS

-----------------------------
3-16-2016
Flushed Zn line back to bubbler 10 times.

Removed manifold from the MOCVD system.
Made sure vacuum chamber was fully sealed to watch if there is a leak.

Dismantled the manifold labeled each valve with its matching number (1~10) which is engraved on the manifold itself.

Cleaned the manifold and valves with toothpicks, IPA and 3 um diamond polish.

Re-connected the valves to the manifold.
Valve #3 and #6 were switched since the #3 valve is till not very clean looking and the #6 valve is not really used because it was for liquid As.
- NS

-----------------------------
3-22-2016
Continued cleaning and vacuum checking manifold.
- NS

-----------------------------
3-23-2016
Re-attached vent line and put the manifold back into the MOCVD system.
Plugged the Fe, As liquid, P liquid lines, replaced pneumatic valves with new ones on the ALPV4 and INPV4. Also removed ALMFC1 S/N 515760588, and install ALMFC1 S/N 021810488.
- NS

-----------------------------
3-24-2016
Baked manifold at 53 degC for ~5.5 hours.
- NS

-----------------------------
3-25-2016
Switched ALMFC1, S/N 021810488 out and put in ALMFC1, S/N 515730588;
Put ALMFC1 S/N 021810488 back in due to a potential vacuum problem with 515730588.
Tape in the TGM was changed.
- NS

-----------------------------
3-28-2016
Flushed PH3 line 5 times;
Flushed TMGa line 5 times;
Flushed AsH3 line 5 times;
Flushed TMIn line 5 times.
- NS

-----------------------------
3-29-2016
Filled PH3 line with PH3;
Filled AsH3 line with AsH3;
Filled TMGa line and TMIn line.
- NS

-----------------------------
3-28-2016
Alarm went off when opening reactor chamber after growth 160329b. Vacuum was not turned off.
- NS

-----------------------------
3-30-2016
Baked manifold over night for ~15 hours at ~58 degC.
- NS

-----------------------------
4-3-2016
Exchanged run-vent connectors
1 <--->  6 front
2 <---> 7
3 <---> 8
4 <--->9
5 <---> 10 (0) back
need to relabel valves.
- PA

-----------------------------
4-6-2016
Noticed vacuum doesn't hold over night.
Test vacuum by plugging run and valve lines in different combinations. Loosened vent-to-manifold connections for lines 6 and 7 and re-tightened them.
- NS

-----------------------------
4-7-2016
Changed TMIn/ TMAl bubbler temp. to 25 degC.
- PA

-----------------------------
4-9-2016
High-temp fault on heater.
restarted, and it appeared again.
May need to increase water flow through heating coil.
- PA

-----------------------------
4-10-2016
Flushed out H2O lines to power supply heat station, tied lines to fan and pump cabinet.
- PA

-----------------------------
4-10-2016
Shaft seems to have shrunk n diameter.
Susceptors are slipping, bore diam too big for shaft. is shaft decomposing on samples?
- PA

-----------------------------
4-12-2016
Replaced front H2 cylinder. New cylinder pressure ~2300 psig.
- NS

-----------------------------
4-20-2016
Steel mesh rolled inside baffle (coffee filter) ~300 um
- PA

-----------------------------
4-20-2016
Flushed TMIn line 5 times.
- NS

-----------------------------
4-20-2016
Test gas flow in reactor chamber with mineral oil smoke.
- NS

-----------------------------
4-20-2016
Changed N2 cylinder. New pressure is 2600 psig.
- NS

-----------------------------
4-23-2016
303 Torr
Cleared out UltraTorr connector.
- PA

-----------------------------
4-28-2016
Baked out manifold over night.
- PA

-----------------------------
4-29-2016
Flushed AsH3 line back to cylinder 5 times;
Flushed PH3 line back to cylinder 5 times;
Flushed In and Ga lines back to bubbler 5 times.
Removed manifold.
- NS

-----------------------------
5-3-2016
Manifold was put back in and put under vacuum.
- NS

-----------------------------
5-4-2016
Tested vacuum with everything hooked up.
Leak rate was ~0.02 Torr/min.
- NS

-----------------------------
5-5-2016
Baked manifold over night. Got to ~60 degC.
- NS

-----------------------------
5-5-2016
TMIn bubbler temp. changed to 17 degC.
- NS

-----------------------------
5-6-2016
37 Torr
- PA

-----------------------------
5-7-2016
641 Torr
Changed pump oil;
Changed oil purifier filters.
- PA

-----------------------------
5-8-2016
120 Torr
- PA

-----------------------------
5-10-2016
New N2 cylinder. ~2300 psig.
- NS

-----------------------------
5-11-2016
702 Torr
- PA

-----------------------------
5-11-2016
Cleaned oven with metal polish.
- NS

-----------------------------
5-12-2016
701 Torr
- PA

-----------------------------
5-12-2016
Switched to non-flexible tube.
- NS

-----------------------------
5-13-2016
350 Torr
- PA

Replaced front H2 cylinder. New cylinder ~2400 psig.
- NS

-----------------------------
5-16-2016
Baked out run line.
Flushed sources.
baking with Al, Ga, In, P, As sources running (purged).

Replaced TMAl MFC
UC2-21S01, S/N 515760588
- PA

-----------------------------
5-17-2016
Evacuated Al, Ga, In, P, As lines 5 times, filled with H2.
- NS

-----------------------------
5-17-2016
Run As and Ga lines with sources open for ~30 min.
- NS

-----------------------------
5-18-2016
Alarm went off when trying to pump down,after idling process.
- NS

-----------------------------
5-19-2016
576.5 Torr
-PA

Smoothed out the fitting on the run line that is in the wall of the MOCVD system.
- NS

-----------------------------
5-27-2016
Replaced front H2 cylinder. ~2400 psig.
- NS

-----------------------------
6-1-2016
Changed N2 cylinder. New pressure 2600 psig.
- NS

-----------------------------
6-2-2016
11.95 Torr
- PA


-----------------------------
6-4-2016
42.7 Torr
- PA

Installed new inlet tube with 1/4 inch UltraTorr on manifold end.
- PA

-----------------------------
6-5-2016
12.7 Torr
 - PA

-----------------------------
6-8-2016
12.2 Torr
- PA

Re-calibrated TMAl MFC S/N 515160588 ,Ultra-Flo.
- PA

-----------------------------
6-9-2016
11.8 Torr
- PA

-----------------------------
6-10-2016
Vented and reinstalled MFC.
- PA

15.6 Torr
Changed to 1/16 inch thermo-couple.
- PA

Flushing TMAl line ~12 times.
- PA

Received new PH3 cylinder.
- PA

-----------------------------
6-11-2016
Replaced gasket on TMAl MFC.
- PA

-----------------------------
6-13-2016
Flushed Al line ~15 times.
- NS

-----------------------------
6-14-2016
Tightened connections on Al line.
- NS

-----------------------------
6-14-2016
Flushed Al line 15 times, finished with N2 carrier gas.
Loosened/ tightened connection between MFC and PV.
Flushed and filled TMAl line with H2 ~5 times.
- NS

-----------------------------
6-16-2016
15.9 Torr
- PA

-----------------------------
6-17-2016
Alarm went off while starting idle procedure.
- NS

-----------------------------
6-22-2016
15.9 Torr
- PA

-----------------------------
6-23-2016
17.8 Torr
- PA

-----------------------------
6-23-2016
11:12 pm
The TGM started to alarm and the TGM read 5 ppb at point #4 system enclosure.
After ~5 seconds, it read 4.9 ppb and the alarm stopped. The reading stayed at 4.9 ppb for ~3.5 min and then read "0".
- NZ

-----------------------------
6-24-2016
12.4 Torr
- PA

-----------------------------
6-27-2016
Changed front H2 cylinder, new pressure ~2300 psig.
- NS

-----------------------------
6-28-2016
Purged TMGa line 3 times, noticed it seemed stable.
- NS

-----------------------------
6-29-2016
11.8 Torr
- PA


-----------------------------
7-3-2016
10.8 Torr
- PA

-----------------------------
7-5-2016
9.9 Torr
- PA

-----------------------------
7-6-2016
11.5 Torr
- PA

Changed N2 cylinder. New pressure ~2400 psig.
- NS

-----------------------------
7-7-2016
11.4 Torr
- PA




Friday, March 11, 2016

Maintenance work of the MOCVD

Historical records will be updated.

For new records, please comment with name, dateitem and reasons. Thanks!

- March-11-2016